# Quantum Bit Threads and the Entropohedron

### v1

@@ 5.6.1 Disentangling lemmas
$\sum_{A\in 2^{X}}\left(\mu^{\prime}(A)+\nu^{\prime}(A)\right)S(A)\leq\sum_{A\in 2^{X}}\left(\mu(A)+\nu(A)\right)S(A)\,.$ (221)
@@
$\sum_{A\in 2^{X}}\left(\mu^{\prime}(A)+\nu^{\prime}(A)\right)S(A)\leq\sum_{A\in 2^{X}}\left(\mu(A)+\nu(A)\right)S(A)\,.$ (225)
@@
Suppose there exists a pair $A,B\in 2^{X}$ with $\mu(A)>0$, $\nu(B)>0$, and $A\cap B\neq\emptyset$. Setting $\alpha:=\min(\mu(A),\nu(B))$, do the following replacement:
@@
leaving $\mu$ and $\nu$ unchanged for all other elements of $2^{X}$. By the WM property (180), this does not increase $\sum_{A}(\mu(A)+\nu(A))S(A)$. It has the following effect on $\phi_{\mu}$, $\phi_{\nu}$:
@@
The proof is similar to the previous one. Suppose there exists a pair of distinct sets $A,B\in 2^{X}$ with $\mu(A)>0$, $\mu(B)>0$, and neither $A\subset B$ nor $B\subset A$. With $\alpha:=\min(\mu(A),\mu(B))$, do the following replacement:
@@
leaving $\mu$ unchanged for all other elements of $2^{X}$. This leaves $\phi_{\mu}$ unchanged, and (by the SSA property (179)) does not increase $\sum_{A}\mu(A)S(A)$.
@@ 5.6.2 Proofs about EDFs
$\text{Maximize }\sum_{i=1}^{m}\sum_{x\in A_{i}}f(x)-\sum_{j=1}^{n}\sum_{x\in B_{j}}f(x)\text{ over $f:X\to\mathbf{R}$ subject to:}\\ \forall\,A\in 2^{X}\,,\quad\left|\sum_{x\in A}f(x)\right|\leq S(A)\,;\qquad\forall\,x\in A_{1}\,,\quad f(x)\geq 0\,;\qquad\forall\,x\in B_{1}\,,\quad f(x)\leq 0\,.$ (234)
@@
$\text{Minimize }\sum_{A\in 2^{X}}\left(\mu(A)+\nu(A)\right)S(A)$ over functions $\mu,\nu:2^{X}\to[0,\infty)$ subject to: (237)
@@
$\text{Minimize }\sum_{i\in{X}^{\prime}}\left|f(x)-f^{\prime}(x)\right|\text{ over $f:X\to\mathbf{R}$ subject to: }\forall\,A\in 2^{X}\,,\quad\left|\sum_{x\in A}f(x)\right|\leq S(A)\,.$ (241)
@@
$\text{Maximize }-\sum_{A\in 2^{X}}\left(\mu(A)+\nu(A)\right)S(A)+\sum_{x\in{X}^{\prime}}(\phi_{\mu}(x)-\phi_{\nu}(x))f^{\prime}(x)\\ \text{ over $\mu,\nu:2^{X}\to[0,\infty)$ subject to: }\\ \forall\,x\in{X}^{\prime}\,,\quad|\phi_{\mu}(x)-\phi_{\nu}(x)|\leq 1\,;\qquad\forall\,x\in{X}\setminus{X}^{\prime}\,,\quad\phi_{\mu}(x)=\phi_{\nu}(x)\,.$ (242)
By corollary 5.18, we can assume that any minimizing pair $\mu$, $\nu$ has disjoint supports. (We will not use the nesting of their supports in this proof, which therefore uses only the WM and not the SSA property of $S$.) The only way to satisfy the second constraint is if $\phi_{\mu}(x)=\phi_{\nu}(x)=0$ for all $x\in{X}\setminus{X}^{\prime}$, in other words $\mu$, $\nu$ are supported only on $2^{{X}^{\prime}}$. The objective then becomes
$\sum_{A\in 2^{{X}^{\prime}}}\left[\mu(A)\left(S(A)-\sum_{x\in A^{\prime}}f^{\prime}(x)\right)+\nu(A)\left(S(A)+\sum_{x\in A^{\prime}}f^{\prime}(x)\right)\right].$ (243)
@@
$\text{Minimize }\sum_{x\in{X}^{\prime}}\left|f(x)-f^{\prime}(x)\right|\text{ over $f:X\to\mathbf{R}$ subject to:}\\ \forall\,A\in 2^{X}\,,\quad\left|\sum_{x\in A}f(x)\right|\leq S(A)\,;\qquad\sum_{x\in{X}\setminus{X}^{\prime}}f(x)=g_{B}\,$, (244)
@@ Acknowledgements.
We would like to thank Juan Pedraza, Andy Svesko, and Brian Swingle for useful discussions. The work of MH is supported by the U.S. Department of Energy through award DE-SC0009986. His also grateful to the Centro de Ciencias de Benasque Pedro Pascual, where part of this work was completed. The work of SRK is supported in part by ISF grant no. 2159/22, by Simons Foundation grant 994296 (Simons Collaboration on Confinement and QCD Strings), by the Minerva foundation with funding from the Federal German Ministry for Education and Research, and by the German Research Foundation through a German-Israeli Project Cooperation (DIP) grant “Holography and the Swampland”. SRK would like to thank Ofer Aharony, Micha Berkooz, Ashoke Sen, and also many people at ICTS Bangalore for useful discussions and for their hospitality during his stay at ICTS. The work of AR is supported by FWO-Vlaanderen project G012222N, the VUB Research Council through the Strategic Research Program High-Energy Physics, and FWO-Vlaanderen through a Senior Postdoctoral Fellowship 1223125N.
@@ Appendix A Comments on generalised entropy
The generalised entropy of a region $r$ is Bousso_2016
@@
That the $\epsilon$-dependence in $S_{\rm gen}$ cancels has only been shown in specific examples Cooperman:2013iqr ; we will assume $S_{\rm gen}$ is always $\epsilon$-independent and finite.

### v2

@@ 5.7.1 Disentangling lemmas
$\sum_{A\subseteq{X}}\left(\mu^{\prime}(A)+\nu^{\prime}(A)\right)S(A)\leq\sum_{A\subseteq{X}}\left(\mu(A)+\nu(A)\right)S(A)\,.$ (223)
@@
$\sum_{A\subseteq{X}}\left(\mu^{\prime}(A)+\nu^{\prime}(A)\right)S(A)\leq\sum_{A\subseteq{X}}\left(\mu(A)+\nu(A)\right)S(A)\,.$ (227)
@@
Suppose there exists a pair of sets $A,B\subseteq X$ with $\mu(A)>0$, $\nu(B)>0$, and $A\cap B\neq\emptyset$. Setting $\alpha:=\min(\mu(A),\nu(B))$, do the following replacement:
@@
leaving $\mu$ and $\nu$ unchanged for all other subsets of X{X}. By the WM property (181), this does not increase $\sum_{A}(\mu(A)+\nu(A))S(A)$. It has the following effect on $\phi_{\mu}$, $\phi_{\nu}$:
@@
The proof is similar to the previous one. Suppose there exists a pair of distinct sets $A,B\subseteq{X}$ with $\mu(A)>0$, $\mu(B)>0$, and neither $A\subset B$ nor $B\subset A$. With $\alpha:=\min(\mu(A),\mu(B))$, do the following replacement:
@@
leaving $\mu$ unchanged for all other subsets of X{X}. This leaves $\phi_{\mu}$ unchanged, and (by the SSA property (180)) does not increase $\sum_{A}\mu(A)S(A)$.
@@ 5.7.2 Proofs about EDFs
$\text{Maximize }\sum_{i=1}^{m}\sum_{x\in A_{i}}f(x)-\sum_{j=1}^{n}\sum_{x\in B_{j}}f(x)\text{ over $f:X\to\mathbf{R}$ subject to:}\\ \forall\,A\subseteq X\,,\quad\left|\sum_{x\in A}f(x)\right|\leq S(A)\,;\qquad\forall\,x\in A_{1}\,,\quad f(x)\geq 0\,;\qquad\forall\,x\in B_{1}\,,\quad f(x)\leq 0\,.$ (236)
@@
$\text{Minimize }\sum_{A\subseteq X}\left(\mu(A)+\nu(A)\right)S(A)$ over functions $\mu,\nu:2^{X}\to[0,\infty)$ subject to: (239)
@@
$\text{Minimize }\sum_{x\in{X}^{\prime}}\left|f(x)-f^{\prime}(x)\right|\text{ over $f:X\to\mathbf{R}$ subject to: }\forall\,A\subseteq X\,,\quad\left|\sum_{x\in A}f(x)\right|\leq S(A)\,.$ (243)
@@
$\text{Maximize }-\sum_{A\subseteq X}\left(\mu(A)+\nu(A)\right)S(A)+\sum_{x\in{X}^{\prime}}(\phi_{\mu}(x)-\phi_{\nu}(x))f^{\prime}(x)\\ \text{ over $\mu,\nu:2^{X}\to[0,\infty)$ subject to: }\\ \forall\,x\in{X}^{\prime}\,,\quad|\phi_{\mu}(x)-\phi_{\nu}(x)|\leq 1\,;\qquad\forall\,x\in{X}\setminus{X}^{\prime}\,,\quad\phi_{\mu}(x)=\phi_{\nu}(x)\,.$ (244)
By corollary 5.18, we can assume that any minimizing pair $\mu$, $\nu$ has disjoint supports. (We will not use the nesting of their supports in this proof, which therefore uses only the WM and not the SSA property of $S$.) The only way to satisfy the second constraint is if $\phi_{\mu}(x)=\phi_{\nu}(x)=0$ for all $x\in{X}\setminus{X}^{\prime}$, in other words $\mu$, $\nu$ are non-zero only on subsets of ${X}^{\prime}$}. The objective then becomes
$\sum_{A\subseteq{{X}^{\prime}}}\left[\mu(A)\left(S(A)-\sum_{x\in A^{\prime}}f^{\prime}(x)\right)+\nu(A)\left(S(A)+\sum_{x\in A^{\prime}}f^{\prime}(x)\right)\right].$ (245)
@@
$\text{Minimize }\sum_{x\in{X}^{\prime}}\left|f(x)-f^{\prime}(x)\right|\text{ over $f:X\to\mathbf{R}$ subject to:}\\ \forall\,A\subseteq X\,,\quad\left|\sum_{x\in A}f(x)\right|\leq S(A)\,;\qquad\sum_{x\in{X}\setminus{X}^{\prime}}f(x)=g_{B}\,$, (246)
@@ Acknowledgements.
We would like to thank Juan Pedraza, Andy Svesko, and Brian Swingle for useful discussions. The work of MH is supported by the U.S. Department of Energy through award DE-SC0009986. He is also grateful to the Centro de Ciencias de Benasque Pedro Pascual, where part of this work was completed. The work of SRK is supported in part by ISF grant no. 2159/22, by Simons Foundation grant 994296 (Simons Collaboration on Confinement and QCD Strings), by the Minerva foundation with funding from the Federal German Ministry for Education and Research, and by the German Research Foundation through a German-Israeli Project Cooperation (DIP) grant “Holography and the Swampland”. SRK would like to thank Ofer Aharony, Micha Berkooz, Ashoke Sen, and also many people at ICTS Bangalore for useful discussions and for their hospitality during his stay at ICTS. The work of AR is supported by FWO-Vlaanderen project G012222N, the VUB Research Council through the Strategic Research Program High-Energy Physics, and FWO-Vlaanderen through a Senior Postdoctoral Fellowship 1223125N.
@@ Appendix A Comments on generalised entropy
The generalised entropy of a region $r$ is Bousso et al. (2016)
@@
That the $\epsilon$-dependence in $S_{\rm gen}$ cancels has only been shown in specific examples Cooperman and Luty (2014); we will assume $S_{\rm gen}$ is always $\epsilon$-independent and finite.